### v3

# Format as a Prior: Quantifying and Analyzing Bias in LLMs for Heterogeneous Data

Abstract
Large Language Models (LLMs) are increasingly employed in applications that require processing information from heterogeneous formats, including texts, tables, infoboxes, and knowledge graphs. However, systematic biases toward particular formats may undermine LLMs’ ability to integrate heterogeneous data impartially, potentially resulting in reasoning errors and increased risks in downstream tasks. Yet it remains unclear whether such biases are systematic, which data-level factors drive them, and what internal mechanisms underlie their emergence.
In this paper, we present the first comprehensive study of format bias in LLMs through a three-stage empirical analysis. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage examines how key data-level factors influence these biases. The third stage analyzes how format bias emerges within LLMs’ attention patterns and evaluates a lightweight intervention to test its effectiveness. Our results show that format bias is consistent across model families, driven by information richness, structure quality, and representation type, and is closely associated with attention imbalance within the LLMs. Based on these investigations, we identify three future research directions to reduce format bias: enhancing data pre-processing through format repair and normalization, introducing inference-time interventions such as attention re-weighting, and developing format-balanced training corpora. These directions will support the design of more robust and fair heterogeneous data processing systems.
Codehttps://github.com/NLPGM/Format-as-a-prior
Appendixhttps://github.com/NLPGM/Format-as-a-prior/appendix.pdf
1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks (Brown et al. 2020). However, their practical deployment remains constrained by key limitations, including factual inaccuracies (commonly referred to as “hallucinations”) (Ji et al. 2023) and incomplete or outdated knowledge (Petroni et al. 2019). One promising direction to address these issues is to incorporate external knowledge sources into the reasoning process—allowing models to ground their outputs in more accurate, up-to-date, and contextually relevant information (Lewis et al. 2020; Gao et al. 2023; Huang and Chang 2022).
In practice, external knowledge exists in diverse formats, ranging from unstructured texts to semi-structured infoboxes, as well as structured tables and Knowledge Graphs (KGs). These different sources of knowledge often complement one another, and the ability to effectively harness them together is crucial for real-world, knowledge-intensive applications.
Figure 1: Format bias affects the LLM’s decision.
The presence of different formats introduces a critical challenge: LLMs may not treat all formats equally when leveraging these heterogeneous data collaboratively. An LLM with strong format preferences interprets information through a distorted lens, giving undue weight to favored formats regardless of actual relevance. This can affect its ability to reason and synthesize effectively.
For example, in clinical decision support, an LLM (e.g., Qwen3-8b) given both textual notes and tabular examination data may overemphasize the text while overlooking key indicators in the table, leading to an incorrect diagnosis. In contrast, as shown in Figure 1, presenting the same information in a knowledge graph enables the model to identify abnormalities and reach the correct conclusion.
When homogeneous inputs are converted into heterogeneous ones with equivalent content, accuracy decreases by 9% and 12% on HotpotQA (Yang et al. 2018) and MuSiQue (Trivedi et al. 2022) (200 samples each), confirming that format heterogeneity can directly impair reasoning performance. This phenomenon may widely arise in heterogeneous reasoning (Christmann et al. 2024), where key evidence is distributed across texts, tables, infoboxes, and knowledge graphs. LLMs may focus on information from their preferred formats, potentially overlooking crucial data in others. Such bias can result in incomplete or flawed conclusions and undermine the LLMs’ role as impartial and effective synthesizers of heterogeneous inputs.
Although there have been several studies exploring various types of bias in LLMs, such as bias between multi-modal data (Zhu et al. 2024; Zhang et al. 2025), there is a lack of systematic research on format bias. To address this gap, we present the first comprehensive investigation and analysis of format bias in LLMs. Our study centers on three critical questions: whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
To systematically investigate the three questions, we conduct a three-stage empirical study by constructing a heterogeneous data conflict scenario for the exploration of bias. The first stage explores the presence and direction of bias across a diverse range of LLMs. The second stage aims to examine how key data-level factors, including information richness, structure quality, and format type, influence these biases. The third stage investigates the emergence of format bias within LLMs’ attention mechanisms and evaluates a lightweight intervention strategy to assess its effectiveness in mitigating such bias.
Our results reveal that format bias is both systematic and consistent across models, driven primarily by differences in information richness, structure quality, and format type. We further show that such bias originates from imbalanced attention allocation during inference and can be partially mitigated through attention-based interventions.
Our key contributions are as follows:
  1. 1.
    To the best of our knowledge, we are the first to investigate the issue of format bias in LLMs and to present a comprehensive investigation of LLM biases toward different knowledge formats across a wide range of LLMs.
  2. 2.
    We conduct a three-stage empirical study to examine the presence and direction of bias, identify the data-level factors that give rise to the bias, and investigate the internal mechanisms in LLMs that contribute to their presence.
  3. 3.
    Based on the comprehensive investigation, we identify three future research directions that may reduce the format bias, which will contribute the development of a more effective heterogeneous data processing system.
2 Related Work
2.1 Heterogeneous reasoning
An important direction in AI research is to develop LLMs capable of reasoning over heterogeneous knowledge sources, including unstructured texts, tables, and KGs, especially when relevant evidence is dispersed across different formats. However, existing methods often struggle to integrate such fragmented information effectively for accurate inference.
To formalize this challenge, recent benchmarks such as COMPMIX (Christmann et al. 2024) and CompMix-IR (Min et al. 2024) require cross-format reasoning, stimulating the development of hybrid QA systems that combine structured and unstructured inputs.
Current approaches can be broadly categorized into two types: (1) Unified retrieval frameworks, which abstract away format heterogeneity using shared APIs or embedding spaces (Xia et al. 2025; Min et al. 2024); and (2) LLM-centric pipelines, which enhance downstream reasoning via evidence selection, re-ranking, or modular tool use (Christmann and Weikum 2024; Lehmann et al. 2024; Zhang et al. 2024; Biswal et al. 2024).
However, existing work often assumes that once evidence is retrieved, LLMs will evaluate it fairly based on content alone. Our study revisits this assumption by asking whether the format in which evidence is presented can influence the model’s judgment, even when the underlying meaning remains the same.
2.2 LLM Behavior under Conflicting Evidence
Recent studies have uncovered a broad range of behavioral biases in LLMs, extending beyond social stereotypes to systematic patterns in reasoning and judgment. A key challenge lies in how LLMs handle conflicts—both between their parametric knowledge (internal beliefs) and in-context evidence, and among competing pieces of evidence (Xu et al. 2024). LLMs tend to favor information aligned with their pre-trained knowledge, even when contradicted by accurate inputs (Jin et al. 2024a; Xie et al. 2023). These LLM biases are shaped by factors such as entity popularity (Xie et al. 2023), event recency (Fang et al. 2023), and evidence frequency (Jin et al. 2024a). Input artifacts also affect LLM behavior, such as preferring self-generated content over retrieved passages (Tan et al. 2024). Related efforts have also examined knowledge conflicts in multi-modal (Zhang et al. 2025; Zhu et al. 2024) and multi-agent settings (Ju et al. 2025), where preference biases and inter-agent inconsistency further complicate reasoning.
Benchmarks like ConflictBank (Su et al. 2024), WikiContradict (Hou et al. 2024), and WhoQA (Pham et al. 2024) have been proposed to evaluate how LLMs handle factual or semantic inconsistencies, especially in ambiguous scenarios. In response, a range of strategies have emerged, including conflict-aware decoding (Yuan et al. 2024; Jin et al. 2024a), counterfactual data augmentation (Fang et al. 2023), internal intervention via attention pruning (Jin et al. 2024b), neuron reweighting (Shi et al. 2024), or prompting LLMs to generate multi-answer responses with source attribution (Shaier et al. 2024).
Our work extends prior research by identifying a previously overlooked source of bias: the format in which information is presented. In contrast to earlier studies that focus on content-level factors such as recency or frequency, we show that differences in format representation alone can systematically influence LLMs’ behavior.
3 Investigation Framework
This section establishes a framework for analyzing format bias in LLMs, encompassing dataset construction, confounding factor exclusion, and automated response evaluation. The dataset construction process is detailed in Appendix A, with evaluation details in Appendix B.
3.1 Dataset and Format Construction
We construct our dataset based on ConflictBank (Su et al. 2024), a public corpus designed to evaluate LLM behavior under factual conflicts. We randomly sample 4,000 entries, each containing one factual claim and three counterclaims with corresponding supporting evidence. This results in 12,000 contradiction pairs, as each claim is paired individually with each of the three counterclaims. Each pair consists of two different claims about the same subject and relation, each supported by its own piece of evidence.
While the original evidence in ConflictBank is presented in plain text, our goal is to examine how LLM behavior is influenced by presenting supporting evidence in different data formats. To construct each heterogeneous contradiction pair, we randomly convert the two pieces of evidence, each of which supports a different claim about the same subject, into different formats. If the selected format is text, no conversion is applied.
Figure 2: Examples of the four data formats used in our experiments: texts, tables, infoboxes, and KGs.
We use GPT-4o-mini as a transformation engine to convert selected texts into one of the following Wikipedia-inspired formats.
  • KGs: A set of (Subject, Predicate, Object) triples capturing core semantic relations.
  • Infobox: A structured key-value format modeled after Wikipedia infobox templates, summarizing factual information.
  • Table: A tabular format styled after Wikipedia tables, presenting comparable facts in labeled rows and columns.
Figure 2 provides illustrative examples of these four formats, showing how semantically equivalent information can be presented in structurally distinct ways. These formats reflect the most commonly used representations in prior work on heterogeneous reasoning. Manual inspection of 5% samples confirms 98.7% factual and 99.3% syntactic accuracy, reflecting strong data fidelity.
3.2 Confounding Factor Control
To ensure that our evaluation isolates the effect of evidence format itself, we implement controls to eliminate two major confounding factors: internal knowledge bias and evidence presentation order.
  • Filtering Internal Knowledge: To ensure that LLM responses are based on external evidence rather than parametric memory, we adopt a filtering procedure consistent with prior work (Gekhman et al. 2024). Each factual claim is tested 16 times by directly querying a given LLM with the corresponding question in a zero-shot setting, and only those samples for which the model fails to reproduce the factual claim in all trials are retained.
  • Randomizing Evidence Order: To eliminate the known bias introduced by input order (Xie et al. 2023), we randomize the sequence of all evidence segments for each input.
3.3 Evaluated LLMs
This evaluation covers ten LLMs across six major series: GPT-4o-mini (Achiam et al. 2023), LLaMA-3.1 (8B), Mistral (7B), Qwen3 (8B, 14B, 30B-A3B, 32B) (Team 2025), Gemma-2 (9B, 27B) (Team et al. 2024), and GLM-4 (9B) (GLM et al. 2024). These LLMs span a range of sizes and architectures, enabling cross-family comparison of format-driven biases.
To ensure reproducibility and eliminate stochasticity, all evaluations were conducted in deterministic inference mode (temperature = 0, without sampling randomness).
3.4 Evaluation Protocol and Metrics
Given the dataset’s scale, we adopt an automated evaluation pipeline using LLMs as adjudicators. Specifically, GPT-4o-mini, GLM-4.5-Air (Zeng et al. 2025), and Qwen-plus are employed to judge which of the two conflicting claims (Source A or Source B) each target response supports.
Each model independently evaluates every sample three times for stability, with the final label per model determined by majority vote. We then compute FPR and DCR for each model and report their averages across the three evaluators. This multi-model setup improves consistency and mitigates variance in individual LLM judgments. Manual checks on a randomly sampled 5% subset show 99.8% agreement between human and averaged LLM judgments, confirming high reliability.
Each LLM response is classified into one of three mutually exclusive categories:
  • Pref-A: The response predominantly or exclusively supports the claim from Source A.
  • Pref-B: The response predominantly or exclusively supports the claim from Source B.
  • Both: The response acknowledges the contradiction and presents both perspectives in a comparative or side-by-side manner.
All responses in our experiments fall unambiguously into one of these three categories, and no additional response types are observed. This categorization enables two quantitative bias metrics used throughout our analysis.:
  • Dual Coverage Rate (DCR): Measures the proportion of responses that acknowledge both claims, indicating an LLM’s capacity to represent multiple perspectives:
    $\text{DCR}=\frac{\text{Both}}{\text{Pref-A}+{\text{Pref-B}+\text{Both}}}$ (1)
  • Format Preference Ratio (FPR): Captures asymmetric bias between conflicting claims when an LLM gives a single-sided response. For the A vs. B experiments, the FPR is calculated as:
    $\text{FPR}=\frac{\text{Pref-A}}{\text{Pref-A}+\text{Pref-B}}$ (2)
These two metrics correspond to the two bias types introduced earlier: DCR captures the presence of bias, while FPR captures the direction of bias.
4 Experimental Results and Analysis
In this section, we present the empirical results of our three-stage investigation. Complete experimental details and raw results are provided in Appendix C.
4.1 Establishing the Existence of Format Bias
Objective and Setup
The primary aim of this initial experiment is to test the null hypothesis that LLMs process information in a format-agnostic manner. To this end, we conduct a large-scale evaluation using the experimental framework outlined in Section 3. We assess ten state-of-the-art LLMs from various families and parameter scales. Each LLM is evaluated on all six possible pairs among the four target data formats: texts, tables, infoboxes, and KGs. This design yields 60 unique experimental conditions (10 LLMs $\times$ 6 format pairs), enabling a comprehensive assessment of systematic format biases.
Top-Level Findings
Figure 3: Average DCR across models under heterogeneous and homogeneous formats. Error bars show standard error.
Our results provide strong evidence against the null hypothesis of format impartiality, revealing instead a consistent and multifaceted pattern of bias in both its presence and direction.
The first pattern, the presence of bias, is pervasive under heterogeneous format conditions. This is reflected in the uniformly low Dual Coverage Rate (DCR), ranging from 3.01% to 24.02%, indicating that LLMs often fail to acknowledge conflicting information across different formats, typically exhibiting a preference for one input while disregarding the other.
To isolate the role of format, we introduce a control condition where both inputs are presented in plain text. As shown in Figure 3, DCR increases markedly under this homogeneous setting. This contrast highlights a broader pattern: format heterogeneity alone can independently and substantially impair a model’s ability to jointly consider multiple inputs, even when the content is semantically equivalent.
This limitation persists across models of varying size, as no clear scaling trend is observed. For example, within the Qwen3 series, larger models do not exhibit improved performance in this regard, suggesting that increased parameter size alone does not resolve this form of processing asymmetry.
Figure 4: Heatmap of FPR between format pairs across LLMs. Asterisks (*) indicate statistical significance under a two-sided binomial test with null hypothesis $\text{FPR}=0.5$.
The second, and more decisive, pattern is a strong directional bias that emerges when an LLM commits to a single-sided response. Despite considerable variation in architecture and scale, LLMs demonstrate a surprisingly consistent pattern of format preferences. As shown in Figure 4, our cross-model analysis reveals a clear preference hierarchy: semantically rich formats such as texts and KGs are consistently favored over visually structured ones like infoboxes and tables. Furthermore, when we group the data by topical domain, we find that these biases persist across domains, suggesting that the observed biases are robust and generalizable. See Appendix C.7 for detailed domain-level results.
4.2 Identifying the Factors Behind Format Bias
Having established the widespread presence of format bias, we now turn to investigating its data-level factors. While prior work has extensively explored factors influencing LLM biases in textual inputs, our focus is on structured data and the properties that may shape LLM behavior in this context. To move from identifying whether such bias exists to understanding why it arises, we decompose the abstract notion of “format” into three representative and controllable dimensions: the structure itself, and the content, which we further divide into quantity and quality.
Building on this decomposition, we hypothesize that three key dimensions (the amount of information conveyed, referred to as quantity; the structural quality of that information, or quality; and the mode of presentation, or format type) play a central role in shaping how LLMs evaluate evidence.
To assess the influence of these factors, we design a unified experimental framework that applies to two of the three factors (with the exception of the format type). Each of these two factors is examined under two complementary conditions:
  • Homogeneous setting: Two evidence sources of the same format are compared, differing only in the factor under investigation. This controlled design isolates the variable to reveal an LLM’s intrinsic sensitivity to that property.
  • Heterogeneous setting: A structured evidence source is paired with unstructured plain texts. This setup assesses the relative strength of the structured format and how its properties influence an LLM’s preference when compared against a universal baseline.
Figure 5: LLM biases across conditions (averaged over ten LLMs). Top: Information Richness; Bottom: Structure Quality. Bars show the proportion of responses favoring the former input, the latter, or both.
Factor 1: Information Richness
This factor concerns the volume of factual detail within an external knowledge source used during reasoning. In real-world applications, the external context provided to LLMs often varies widely in the amount of information it contains. It serves as a proxy for the completeness of factual detail. To examine whether LLMs apply a “more is better” heuristic, interpreting quantity as indicative of evidentiary strength, we systematically vary the number of entries in structured formats (e.g., table rows or knowledge graph triples).
In the homogeneous setting, we conduct experiments within each format type (tables, KGs, and infoboxes), comparing three levels of information richness across three pairwise conditions: 4 vs. 8 entries, 8 vs. 12 entries, and 4 vs. 12 entries.Results consistently indicate that LLMs favor the richer variant in each pair, irrespective of format. This suggests that even when structure is held constant, LLMs exhibit a systematic bias for inputs with more factual content (see Figure 6).
In the heterogeneous setting, we assess whether this preference generalizes when structured inputs are compared against unstructured texts. For each format, we construct three pairs: texts vs. 4-entry structure, texts vs. 8-entry structure, and texts vs. 12-entry structure. All structured inputs are generated from the same source texts as the corresponding text versions, ensuring content consistency. Across all formats, LLMs’ bias for the structured inputs increases with the number of entries (see Figure 6).
These findings suggest that LLMs tend to associate greater volumes of factual content with higher evidentiary value, both within individual formats and when comparing structured and unstructured inputs.
Figure 6: Average FPR for structured data vs. texts across ten LLMs. Top: Information Richness; Bottom: Structure Quality. Shaded areas indicate mean $\pm$ one standard deviation.
Factor 2: Structure Quality
This factor examines whether LLMs are sensitive to the structural integrity of external knowledge inputs. In practice, structured formats like tables or knowledge graphs may contain noise or malformed syntax. To simulate this, we introduce controlled corruption into structure-defining tokens (e.g., brackets, colons, separators), randomly replacing them with other characters or blanks at fixed probabilities (0.45 and 0.9), while preserving the underlying factual content.
In the homogeneous setting, we compare clean and corrupted versions within each format type. LLMs consistently favor the well-formed input, confirming that structure quality serves as a reliability signal. Notably, the preference saturates beyond moderate corruption (e.g., 0.45), indicating that LLMs tend to treat inputs as either structurally valid or invalid (see Figure 6).
In the heterogeneous setting, we pair each corrupted version with its corresponding clean texts. Each corruption level is applied to the same intact structured data instance, guaranteeing that the underlying information remains identical across differently corrupted versions. LLMs’ bias for the structured inputs declines sharply as corruption increases, despite identical semantics. This suggests that structural degradation alone can undermine the perceived credibility of otherwise accurate structured inputs (see Figure 6).
Factor 3: Format Type
This factor investigates the impact of the representational structure and layout used to represent logically equivalent information. The choice between plain text, a relational graph (KGs), or a visual grid (tables, infoboxes) embodies core differences in format semantics. Holding the content constant, we ask: do LLMs possess intrinsic preferences for certain data structures?
Metric Infobox Table KGs
FPR 0.235 0.398 0.336
Table 1: Average FPR between structured data and texts (mean across ten LLMs).
We construct matched pairs with identical factual entries in both texts and structured variants. As shown in Table 1, the results reveal a consistent hierarchy: tables are most competitive, followed by KGs, with infoboxes least preferred. All three structured variants contain nearly identical content, differing only in their representational layout. This suggests that the format itself, rather than the informational content, modulates LLMs’ bias.
Cross-Factor Insight
Consistent with the findings in Section 4.1, we observe that format homogeneity substantially reduces bias in presence. This trend holds not only for unstructured inputs but also extends to structured formats such as tables and knowledge graphs, where DCR increases significantly (28%–53%) when both inputs adopt the same format. These results suggest that LLMs are more capable of jointly considering multiple sources of information when they are presented in a uniform structure. In contrast, even when the content is semantically equivalent, presenting information in heterogeneous formats tends to impair integration and leads to partial or selective processing.
4.3 Mechanism Behind Format Bias
In this analysis, we move beyond identifying data-level factors to analyzing how format bias manifests within the internal processing of LLMs. Our aim is to examine whether differences in attention allocation across input formats are associated with the presence and direction of bias identified earlier. We focus our analysis on three representative LLMs from different model families: Qwen3-8B, Mistral-7B-Instruct-v0.3, and Llama-3.1-8B-Instruct.
The Relation Between Attention Allocation and Presence of Bias
We begin by analyzing how attention is distributed between conflicting evidence inputs during inference. For each input pair, we compute the mean attention mass assigned to each segment and then calculate the absolute difference between the two values to quantify the degree of imbalance.
To quantify the negative correlation between attention gap and DCR, we employ Spearman’s rank correlation coefficient as a measurement. The results show coefficients of –0.31, –0.37, and –0.54 across the three LLMs, indicating a weak to moderate negative correlation. This suggests that greater imbalance in attention distribution is associated with a lower likelihood of the model recognizing both sources of information.
This implies that format bias arises, at least in part, from skewed attention allocation early in processing, which increases the chance of one source being overlooked.
The Relation Between Attention Allocation and Direction of Bias
We further examine whether attention allocation can explain which input LLMs tend to prefer when selecting only one side. Interestingly, in 82.35 percent of such cases, they favor the source that received less attention.
This observation implies that while attention imbalance affects whether both sources are represented in the output, it does not consistently determine the direction of preference. In other words, a segment receiving more attention does not necessarily have a higher chance of being selected as the final answer.
These results indicate that attention allocation is related to both types of bias, though the nature of this relationship may differ.
Figure 7: Effects of intervention methods in terms of DCR and FPR.
Attention-Guided Intervention
To move from correlation to causation, we design an intervention that directly modifies the internal representations produced by the attention mechanism.
Specifically, we apply a normalization-based reweighting to the attention distribution at each generation step. Let $A\in\mathbb{R}^{H\times L_{q}\times L_{k}}$ denote the attention tensor, where $H$ is the number of heads, $L_{q}$ the query sequence length, and $L_{k}$ the key sequence length. At the current decoding step, we denote the attention distribution over keys as $a\in\mathbb{R}^{L_{k}}$, where $a_{j}$ represents the attention weight assigned to the $j$-th key token.
To ensure that the two pieces of evidence receive equal total attention, we define the corresponding token index sets $I_{1}$ and $I_{2}$, and compute the total attention mass on each:
$m_{1}=\sum_{j\in I_{1}}a_{j},\quad m_{2}=\sum_{j\in I_{2}}a_{j}$ (3)
We then compute their average:
$\bar{m}=\frac{m_{1}+m_{2}+\varepsilon}{2}$ (4)
where $\varepsilon$ is a small constant for numerical stability. The attention weights are reweighted accordingly:
$a^{\prime}_{j}=\begin{cases}\frac{\bar{m}}{m_{1}+\varepsilon}\cdot a_{j},&j\in I_{1}\\ \frac{\bar{m}}{m_{2}+\varepsilon}\cdot a_{j},&j\in I_{2}\\ a_{j},&\text{otherwise}\end{cases}$ (5)
This procedure enforces equal total attention mass across the two evidence segments, while preserving the original intra-segment distribution.
Experiment Results
We apply the attention-balancing intervention to three representative LLMs and compare their behavior before and after modification.
The results reveal a consistent improvement in the models’ ability to attend to both conflicting sources: DCR significantly increases across all format pairs and all three models (see Figure 7). This suggests that enforcing a more balanced attention distribution during inference effectively encourages the model to integrate information from both input segments, rather than overlooking one entirely. When applied to heterogeneous-input reasoning tasks, the attention-balancing method also improves performance on downstream RAG-style QA datasets, including HotpotQA and MuSiQue, where accuracy increases by 6.5% and 9.5%, respectively.
In contrast, FPR remains largely stable after the intervention, and the changes are not statistically significant across the evaluated models (see Figure 7). This implies that although the LLMs process both sources more evenly, the intervention has limited effect on their final output preferences once a directional bias has emerged.
The different effects of attention interventions on DCR and FPR suggest that presence of biases are at least partially controllable at inference time, whereas direction of biases are more resistant to modification. These more stable preferences may reflect deeper inductive biases acquired during pretraining, such as stronger alignment with textual inputs.
5 Discussion
Our findings show that format bias in LLMs is systematic and has practical consequences for systems processing heterogeneous data. To mitigate it, interventions can be applied at three levels: pre-processing, inference, and model development.
Data Pre-processing
Pre-processing is a cost-effective way to reduce bias. Since LLMs tend to trust well-structured inputs, automatically repairing corrupted format in tables or KGs can prevent valuable content from being dismissed. Additionally, using a consistent input format can help reduce format-induced bias, as LLMs perform better when differences in input format is minimized. During the transformation process, it is important to preserve all essential information while avoiding the introduction of noise or unintended alterations.
Inference-Time Intervention
Re-balancing attention across inputs during inference can help LLMs more effectively integrate information from multiple sources, reducing the tendency to overlook less-preferred formats. By encouraging the model to distribute attention more evenly across heterogeneous data, it enhances the model’s ability to incorporate information from all inputs and supports more robust and balanced reasoning. Although such adjustments may not fully shift the model’s final output bias, they improve its intermediate processing. Future work may explore more fine-grained or deeper intervention strategies to better align model attention with content relevance and reduce structural bias during inference.
Model Development and Fine-tuning
Format preferences likely originate from pretraining data imbalance. Mitigation may involve training on format-balanced corpora, using contrastive objectives, or incorporating format-aware modules. Though costlier than inference-time methods, such approaches offer a more fundamental solution.
6 Conclusion
This work presents the first in-depth study of format bias in LLMs, identifying it as a consistent effect driven by information richness, structure quality, and format type. The bias manifests in two types: the presence of bias that can be mitigated, and the direction of bias likely rooted in pretraining. These findings underscore the importance of treating data format as a core factor in LLM design and evaluation. We also outline practical mitigation strategies, including data preprocessing, inference-time attention adjustment, and format-aware training, which together offer clear paths for reducing format bias in heterogeneous reasoning.
Acknowledgments
We sincerely thank the anonymous reviewers for their insightful comments and helpful suggestions, which helped improve this work.
This work was supported by the National Natural Science Foundation of China (NSFC) (Grant No. 62576256) and the Fundamental Research Funds for the Central Universities, China (Grant No. 2042022dx0001).
Appendix
Appendix A Format Conversion Details
A.1 Prompt Templates
We provide two types of prompt templates used to convert unstructured text into structured formats. The first type does not specify the number of resulting entries, allowing models to extract as many relevant facts as they find appropriate. This type is used to evaluate the model’s natural preference when unconstrained. The second type explicitly constrains the number of facts to be included, ensuring consistency in information quantity across formats and facilitating controlled experiments on information richness.
For each type, we provide templates targeting three structured formats: table, knowledge graph (KG), and infobox, enabling structured data generation in different representation styles while maintaining semantic equivalence with the original text.
Type I: Format Conversion without Entry Constraints
Text to Table
Prompt is as follows:
1## ROLE & GOAL
2You are an expert data architect. Your goal is to convert the provided [Source Text] into a structured, clear, and accurate **Wikipedia-style MediaWiki table**.
3
4## CRITICAL RULE
5**RULE #1: The [Claim to Prioritize] MUST be perfectly and centrally represented in the table. This claim is the most important piece of information.**
6
7## RULES
82. Begin the table with ‘{{| class="wikitable"‘ and end with ‘|}}‘.
93. Add a descriptive caption using ‘|+ Caption text (summarize the table purpose clearly).
104. Choose column headers that clearly categorize the core facts.
115. Only include rows and columns directly supported by the [Source Text]. Do not infer or add information.
126. Use ‘!‘ for headers and ‘|‘ or ‘||‘ for data cells. Use ‘|-‘ to separate rows.
137. Ensure formatting is clean and valid per MediaWiki syntax.
14
15## EXAMPLE
16- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
17- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
18- **Output Table:**
19{{| class="wikitable"
20|+ Details of Grave of the Fireflies
21|-
22! Film Title !! Director !! Studio !! Runtime
23|-
24| Grave of the Fireflies $|$ Isao Takahata $|$ Studio Ghibli $|$ 89 minutes
25|}}
26
27---
28## YOUR TASK
29
30[Claim to Prioritize]:
31{claim_text}
32
33[Source Text]:
34{evidence_text}
35
36[Output MediaWiki Table]:
Text to Knowledge Graph (KG)
Prompt is as follows:
1## ROLE & GOAL
2You are a knowledge engineer. Your goal is to extract all factual relationships from the [Source Text] and represent them as (Subject, Predicate, Object) triplets.
3
4## CRITICAL RULE
5**RULE #1: The [Claim to Prioritize] MUST be converted into one or more primary, accurate triplets. This claim is the most important piece of information.**
6
7## RULES
82. Each triplet must be on a new line and enclosed in parentheses ‘()‘.
93. All triplets must be directly derivable from the [Source Text]. Do not make assumptions.
104. Use consistent and clear names for entities and predicates.
115. Extract one triplet for each distinct fact.
12
13## EXAMPLE
14- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
15- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
16- **Output Triplets:**
17(Grave of the Fireflies, has_director, Isao Takahata)
18(Grave of the Fireflies, has_studio, Studio Ghibli)
19(Grave of the Fireflies, has_runtime_minutes, 89)
20(Grave of the Fireflies, release_year, 1988)
21
22---
23## YOUR TASK
24
25[Claim to Prioritize]:
26{claim_text}
27
28[Source Text]:
29{evidence_text}
30
31[Output Triplets]:
Text to Infobox
Prompt is as follows:
1## ROLE & GOAL
2\#You are a meticulous Wikipedia editor. Your goal is to summarize the key facts from the [Source Text] into a concise infobox format.
3
4\### CRITICAL RULE
5\#**RULE #1: The [Claim to Prioritize] MUST be accurately included as a key-value pair in the infobox. This claim is the most important piece of information.**
6
7\### RULES
8\#2. The format must follow the MediaWiki infobox style, like:
9\#{{{{Infobox [type]
10\#| key1 = value1
11\#| key2 = value2
12\#...
13\#}}}}
14\#3. Use a relevant infobox type in the first line (e.g., book‘, film‘, person‘, etc.), based on the source content.
15\#4. Only include information explicitly mentioned in the [Source Text].
16\#5. Field names (keys) should be relevant, standard when possible, but flexible based on content.
17\#6. Values should be brief and precise. No full sentences.
18\#7. Do not add, infer, or assume any details not supported by the source.
19
20\### EXAMPLE
21\#- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
22\#- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
23\#- **Output Infobox:**
24\#{{{{Infobox film
25\#| title = Grave of the Fireflies
26\#| director = Isao Takahata
27\#| studio = Studio Ghibli
28\#| runtime = 89 minutes
29\#| year = 1988
30\#}}}}
31
32\#---
33\### YOUR TASK
34
35\#[Claim to Prioritize]:
36\#{claim_text}
37
38\#[Source Text]:
39\#{evidence_text}
40
41\#[Output Infobox]:
Type II: Format Conversion with Entry Constraints
Text to Table with <nums> Facts
Prompt is as follows:
1## ROLE & GOAL
2You are an expert data architect. Your goal is to convert the provided [Source Text] into a structured, clear, and accurate **Wikipedia-style MediaWiki table**, containing **exactly {nums} key facts**.
3
4## CRITICAL RULE
5**RULE #1: The [Claim to Prioritize] MUST be perfectly and centrally represented in the table. This claim is the most important piece of information.**
6
7## RULES
82. Begin the table with ‘{{| class="wikitable"‘ and end with ‘|}}‘.
93. Add a descriptive caption using ‘|+ Caption text (summarize the table purpose clearly).
104. Choose column headers that clearly categorize the core facts.
115. Only include rows and columns directly supported by the [Source Text]. Do not infer or add information.
126. Use ‘!‘ for headers and ‘|‘ or ‘||‘ for data cells. Use ‘|-‘ to separate rows.
137. Include **exactly {nums} distinct facts** across the table.
148. Ensure formatting is clean and valid per MediaWiki syntax.
15
16## EXAMPLE
17- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
18- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
19- Output MediaWiki Table with Exactly 4 Facts
20- **Output Table:**
21{{| class="wikitable"
22|+ Details of Grave of the Fireflies
23|-
24! Film Title !! Director !! Studio !! Runtime
25|-
26| Grave of the Fireflies $|$ Isao Takahata $|$ Studio Ghibli $|$ 89 minutes
27|}}
28
29---
30
31## YOUR TASK
32
33[Claim to Prioritize]:
34{claim_text}
35
36[Source Text]:
37{evidence_text}
38
39[Output MediaWiki Table with Exactly {nums} Facts]:
Text to Knowledge Graph with <nums> Triples
Prompt is as follows:
1## ROLE & GOAL
2You are a knowledge engineer. Your goal is to extract **exactly {nums}** factual relationships from the [Source Text] and represent them as (Subject, Predicate, Object) triplets.
3
4## CRITICAL RULE
5RULE #1: The [Claim to Prioritize] MUST be converted into one or more triplets within the {nums} total. This claim is the **top priority** and must be captured accurately.
6
7## EXTRACTION RULES
82. You must extract **exactly {nums}** triplets. No more, no fewer.
93. Each triplet must be on a new line and enclosed in parentheses ‘()‘.
104. All triplets must be directly supported by the [Source Text]. Do **not** infer or assume anything not explicitly stated.
115. Use clear and consistent naming for subjects, predicates, and objects.
126. Each triplet must capture a distinct fact.
13
14## EXAMPLE
15- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
16- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
17- **Output Triplets:**
18(Grave of the Fireflies, has_producer, Studio Ghibli)
19
20---
21
22## YOUR TASK
23
24[Claim to Prioritize]:
25{claim_text}
26
27[Source Text]:
28{evidence_text}
29
30[Output Triplets with Exactly {nums} Triplets]:
Text to Infobox with <nums> Key-Value Pairs
Prompt is as follows:
1## ROLE & GOAL
2You are a meticulous Wikipedia editor. Your goal is to summarize the key facts from the [Source Text] into a concise infobox format, with **exactly {nums} key-value pairs**.
3
4## CRITICAL RULE
5**RULE #1: The [Claim to Prioritize] MUST be accurately included as a key-value pair in the infobox. This claim is the most important piece of information.**
6
7## RULES
82. The format must follow the MediaWiki infobox style, like:
9{{{{Infobox [type]
10| key1 = value1
11| key2 = value2
12...
13}}}}
143. Use a relevant infobox type in the first line (e.g., book‘, film‘, person‘, etc.), based on the source content.
154. Only include information explicitly mentioned in the [Source Text].
165. Field names (keys) should be relevant, standard when possible, but flexible based on content.
176. Values should be brief and precise. No full sentences.
187. Include **exactly {nums} key-value pairs**. No more, no fewer.
198. Do not add, infer, or assume any details not supported by the source.
20
21## EXAMPLE
22- **Source Text:** "The 1988 film Grave of the Fireflies was directed by Isao Takahata and produced by Studio Ghibli. Its runtime is 89 minutes."
23- **Claim to Prioritize:** "’Grave of the Fireflies was produced by Studio Ghibli."
24- Output Infobox with Exactly 5 Key-Value Pairs.
25- **Output Infobox:**
26{{{{Infobox film
27| title = Grave of the Fireflies
28| director = Isao Takahata
29| studio = Studio Ghibli
30| runtime = 89 minutes
31| year = 1988
32}}}}
33
34---
35
36## YOUR TASK
37
38[Claim to Prioritize]:
39{claim_text}
40
41[Source Text]:
42{evidence_text}
43
44[Output Infobox with Exactly {nums} Key-Value Pairs]:
A.2 Examples of Format Transformation
We provide representative examples of how the same source text is converted into different structured formats. Each example includes: (1) the original unstructured input, (2) the core claim to be emphasized during conversion, and (3) the transformed output in the target format.
Example 1: Text to Table
Source Text
As the sun dipped into the Mediterranean, casting a warm orange glow over the bustling streets of Barcelona, Marie-France de Rose made her way through the crowded corridors of the Provincial Council building. Her heels clicked against the polished marble floor, a rhythmic accompaniment to the hum of conversation and the rustle of papers being shuffled. She exchanged warm smiles and nods with colleagues and staff members, her bright blue eyes sparkling with a sense of purpose. As she approached the door to her office, a young intern, Jordi, hurried to open it for her. D̈eputy de Rose, you have a meeting with the mayor in fifteen minutes,ḧe reminded her, his eyes darting to the stack of files in her arms. Marie-France nodded, her dark hair bobbing with the movement. G̈ràcies, Jordi. I’ll review the proposals one more time before the meeting.S̈he slipped into her office, the scent of fresh coffee and citrus wafting from the cup on her desk. As she began to sift through the documents, her gaze lingered on the crest of the Provincial Council emblazoned on the letterhead: a stylized image of Saint George, the patron saint of Catalonia, surrounded by the words D̈iputació de Barcelonaïn bold, modern script. It was a symbol of her pride and responsibility as a Barcelona provincial deputy, a role she had worked tirelessly to earn. The sound of her phone buzzing broke her concentration. She glanced at the screen, her eyes narrowing as she read the message from her colleague, Xavier. M̈eet me in the courtyard at 5. We need to discuss the latest developments on the La Rambla revitalization project.M̈arie-France’s lips curved into a determined smile. This was exactly the kind of initiative she had been championing as deputy – projects that would breathe new life into the city’s historic heart and benefit its citizens. With a sense of anticipation, she tucked the phone into her pocket and returned to her preparations for the meeting with the mayor.
Claim to Prioritize
Marie-France de Rose holds the position of Barcelona provincial deputy.
Converted Output (Table)
The converted result is as follows:
[…]
Example 2: Text to Knowledge Graph (KG)
Source Text
**Breaking News: Marie-France de Rose Confirmed as Member of European Parliament** In a move that solidifies her position as a prominent figure in European politics, Marie-France de Rose has been officially recognized as a member of the European Parliament. This esteemed title is a testament to her dedication to serving the people of France and the European Union as a whole. According to the European Parliament’s official website, Marie-France de Rose is listed as a member of the Parliament, representing the interests of French citizens since her election. This confirmation is further supported by multiple credible sources, including the French Ministry of Foreign Affairs and the European Commission. M̈arie-France de Rose has been a vocal advocate for European integration and has worked tirelessly to promote the interests of France and the EU,s̈aid Jean-Luc Mélenchon, a fellow French politician and leader of the La France Insoumise party. Ḧer election to the European Parliament is a well-deserved recognition of her commitment to public service.D̈ata from the European Parliament’s election results in 2019 shows that Marie-France de Rose received a significant number of votes, securing her position as a member of the Parliament. According to the official results, she garnered over 1.2 million votes, making her one of the top vote-getters in the French delegation. M̈arie-France de Rose’s election to the European Parliament is a testament to her hard work and dedication to the people of France,s̈aid Nathalie Loiseau, a French politician and former Minister for European Affairs. S̈he has been a strong voice for French interests in the Parliament and has worked to build bridges between France and other EU member states.Äs a member of the European Parliament, Marie-France de Rose will continue to play a crucial role in shaping EU policy and legislation, working closely with other MEPs to address the pressing issues facing the European Union. In conclusion, the evidence overwhelmingly supports the claim that Marie-France de Rose holds the position of member of the European Parliament.
Claim to Prioritize
Marie-France de Rose holds the position of member of the European Parliament.
Converted Output (KG Triples)
The converted result is as follows:
1"(Marie-France de Rose, holds_position, member of the European Parliament)
2(Marie-France de Rose, is_recognized_as, member of the European Parliament)
3(Marie-France de Rose, represents, interests of French citizens)
4(Marie-France de Rose, elected_in, 2019)
5(Marie-France de Rose, received_votes, over 1.2 million)
6(Marie-France de Rose, is_top_vote_getter, in French delegation)
7(Marie-France de Rose, is_advocate_for, European integration)
8(Marie-France de Rose, works_to_promote, interests of France and the EU)
9(Marie-France de Rose, plays_role, in shaping EU policy and legislation)
10(Marie-France de Rose, works_with, other MEPs)
Example 3: Text to Infobox
Source Text
**BARCELONA, SPAIN - MARCH 15, 2042** After Seven Years of Dedicated Service, Marie-France de Rose Steps Down as Barcelona Provincial Deputy In a move that marks the end of an era, Marie-France de Rose, the French politician who has been a stalwart of the Barcelona administration, has stepped down from her position as Barcelona provincial deputy, a role she has held since 2035. During her seven-year tenure, de Rose has been instrumental in shaping the city’s policies and initiatives, earning her widespread respect and admiration from her peers and constituents alike. Her dedication to the people of Barcelona has been unwavering, and her commitment to improving the quality of life for all citizens has been a hallmark of her time in office. Ï am proud of what we have achieved together,d̈e Rose said in an exclusive interview with our newspaper. F̈rom implementing sustainable transportation systems to promoting cultural exchange programs, we have worked tirelessly to make Barcelona a better place for everyone. It’s been an honor to serve this great city, and I am confident that the progress we’ve made will continue to benefit future generations.Ünder de Rose’s leadership, Barcelona has seen significant improvements in areas such as education, healthcare, and environmental sustainability. Her administration’s efforts to reduce carbon emissions and increase the use of renewable energy sources have earned the city international recognition and praise. M̈arie-France de Rose has been a true champion of Barcelona,s̈aid Mayor Xavier Domènech, who worked closely with de Rose during her tenure. Ḧer vision, expertise, and passion for the city have inspired us all, and her legacy will continue to inspire future leaders. We will miss her dedication and commitment, but we are grateful for the time she has given to our city.D̈e Rose’s decision to step down has sparked speculation about her future plans, with many wondering if she will return to French politics or pursue other opportunities. When asked about her next move, de Rose remained tight-lipped, saying only that she was looking forward to taking a well-deserved break and spending time with her family. As Barcelona bids farewell to a dedicated public servant, the city’s residents and officials alike are grateful for the impact Marie-France de Rose has had on their lives.”
Claim to Prioritize
Marie-France de Rose holds the position of Barcelona provincial deputy form 2035 to 2042.
Converted Output (Infobox)
The converted result is as follows:
1"{{Infobox person
2 | name = Marie-France de Rose
3 | position = Barcelona provincial deputy
4 | term_start = 2035
5 | term_end = 2042
6 | nationality = French
7 | notable_work = Policies and initiatives in Barcelona
8}}"
Following prior work (Min et al. 2024; Biswal et al. 2024; Christmann and Weikum 2024; Xia et al. 2025), we directly feed the transformed structured data into the LLMs without additional adapters or format-specific encoders.
A.3 Data Verification Statistics
To ensure the integrity and reliability of our format transformation pipeline, we conducted a manual verification over a randomly sampled subset of 600 examples (5% of the full dataset). Each sample was independently assessed according to two criteria:
  • Factual Consistency: Whether the transformed format preserves the key factual content from the original text without hallucination or distortion.
  • Syntax Validity: Whether the output adheres to the syntactic conventions of the target format (e.g., MediaWiki table syntax, well-formed triples, valid infobox fields).
We observed high fidelity across both dimensions, with 592 out of 600 samples passing the factual consistency check and 596 passing the syntax validity check. These results confirm that the conversion process produces high-quality structured representations suitable for downstream analysis.
Appendix B Model Query and Evaluation Protocols
B.1 Evaluated LLMs
  • Qwen3-8B-FP8
    HuggingFace: Qwen/Qwen3-8B-FP8
  • Qwen3-14B-FP8
    HuggingFace: Qwen/Qwen3-14B-FP8
  • Qwen3-32B-FP8
    HuggingFace: Qwen/Qwen3-32B-FP8
  • Qwen3-30B-A3B-FP8
    HuggingFace: Qwen/Qwen3-30B-A3B-FP8
  • Llama-3.1-8B-Instruct-FP8
    HuggingFace: nvidia/Llama-3.1-8B-Instruct-FP8
  • gemma-2-9b-it-FP8
    HuggingFace: RedHatAI/gemma-2-9b-it-FP8
  • gemma-2-27b-it-FP8
    HuggingFace: nm-testing/gemma-2-27b-it-FP8
  • glm-4-9b-chat-hf
    HuggingFace: zai-org/glm-4-9b-chat-hf
  • Mistral-7B-Instruct-v0.3
    HuggingFace: mistralai/Mistral-7B-Instruct-v0.3
  • Llama-3.1-8B-Instruct
    HuggingFace: meta-llama/Llama-3.1-8B-Instruct
  • Qwen3-8B
    HuggingFace: Qwen/Qwen3-8B
B.2 Decoding and Evaluation Parameters
  • Decoding temperature: 0
  • Evaluation method: GPT-4o-mini, 3-pass majority voting
B.3 Answer Prompt Templates
We provide two distinct prompting templates used to elicit answers from LLMs under different conditions: (1) without any supporting evidence to test internal (parametric) knowledge, and (2) with explicit reference sources to simulate evidence-based reasoning.
Prompt 1: Answer Without Reference (Internal Knowledge Only)
This prompt is used in the internal knowledge filtering phase (see Section 3.2). It encourages the model to produce a concise, self-contained answer without relying on any external input.
1Answer the question with a single word or phrase. Do not explain or add any other content.
Prompt 2: Answer With Reference (Evidence-Based Inference)
This prompt is used in the main evaluation setup where the model is required to synthesize information from two conflicting evidence sources. The template injects both sources as contextual grounding and instructs the model to answer concisely.
1Based on the two reference sources provided below, answer the following question **concisely**.
2
3{full_reference}
4
5Question: {question}
B.4 LLM Evaluator Prompt
1Question:
2{question}
3Answer:
4{answer}
5Claim A:
6{claim_shared}
7Claim B:
8{claim_specific}
9Task:
10Evaluate how closely the answer aligns with Claim A and Claim B using the scale below.
11Scoring Guidelines:
121 - The answer aligns only with Claim A
132 - The answer aligns with both Claim A and Claim B
143 - The answer aligns only with Claim B
15No - The answer aligns with neither claim
16Provide only a single score: 1, 2, 3, or No.
To assess the reliability of the automatic evaluation, we manually annotated a random 5% subset of the evaluation data, totaling 600 examples. Among these, 599 model judgments matched the human annotations, demonstrating a high degree of alignment and validating the accuracy of the automated scoring process.
Appendix C Full Results
C.1 Response Counts by Format Pairing
To provide a complete view of model behavior across different format conditions, we report the raw count of response types: Pref-A, Pref-B, and Both, for each model and each of the six heterogeneous format pairings. Each entry corresponds to the number of samples in which the model supported the first format (Pref-A), the second format (Pref-B), or acknowledged both (Both). These results form the basis for the DCR and FPR metrics used in Section 4. Full results are shown in Table 78.
In addition, we report the response distributions for a control condition where both conflicting inputs are presented in plain text. This setting allows us to isolate the effect of format heterogeneity by establishing a baseline for dual coverage when content format is held constant. Table 2 summarizes, for each model, the number of cases where Both claims were acknowledged and the total number of evaluated examples. This comparison highlights the substantial drop in dual-claim recognition under heterogeneous conditions, emphasizing the impact of format presentation on information integration.
Model Both Total
Llama-3.1-8B-Instruct-FP8 3154 10035
gpt4omini 4645 11032
Qwen3-8B-FP8 3120 10225
Qwen3-14B-FP8 4008 11785
Qwen3-32B-FP8 3479 11759
Qwen3-30B-A3B-FP8 3514 11012
gemma-2-9b-it-FP8 2656 11797
gemma-2-27b-it-FP8 5190 11836
glm-4-9b-chat-hf 1256 11159
Mistral-7B-Instruct-v0.3 3120 11575
Table 2: Dual Coverage in Text-Only Control Condition
C.2 Effect of Information Richness
We examine the impact of information richness on LLM preferences by varying the number of factual entries presented in each structured format. Table 910 presents results under homogeneous format settings, where structured inputs with different levels of richness are compared. Table 1314 shows the results under heterogeneous settings, where structured inputs are compared against plain text versions with equivalent content. These tables allow us to quantify the degree to which factual quantity influences model preferences.
C.3 Effect of Structure Quality
We assess how the structural quality of inputs affects model behavior by introducing controlled corruption into format-specific tokens. Table 1112 presents results under homogeneous settings, where clean and corrupted structured inputs are compared. Table 1516 shows the results under heterogeneous settings, where corrupted structured inputs are compared against clean text. These results highlight the sensitivity of LLMs to syntactic well-formedness.
C.4 Effect of Format Type
We investigate whether models exhibit consistent preferences among different format types when content is held constant. Table 3 presents results comparing plain text with each of the three structured alternatives: tables, infoboxes, and knowledge graphs. This experiment isolates the influence of format representation on model preference.
Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
infobox vs texts 487 1170 186 1843
tables vs texts 725 927 195 1847
kg vs texts 556 1208 70 1834
GPT-4o-mini
infobox vs texts 333 1353 179 1865
tables vs texts 532 907 432 1871
kg vs texts 503 1030 291 1824
Qwen3-8B-FP8
infobox vs texts 439 1413 73 1925
tables vs texts 692 1028 205 1925
kg vs texts 630 1011 138 1779
Qwen3-14B-FP8
infobox vs texts 447 1293 177 1917
tables vs texts 580 958 383 1921
kg vs texts 583 1123 211 1917
Qwen3-32B-FP8
infobox vs texts 398 1371 132 1901
tables vs texts 683 890 331 1904
kg vs texts 650 1015 237 1902
Qwen3-30B-A3B-FP8
infobox vs texts 369 1445 94 1908
tables vs texts 651 1002 265 1918
kg vs texts 627 1118 166 1911
Gemma-2-9b-it-FP8
infobox vs texts 521 1329 84 1934
tables vs texts 657 1123 152 1932
kg vs texts 490 1340 99 1929
Gemma-2-27b-it-FP8
infobox vs texts 513 1306 161 1980
tables vs texts 782 844 349 1975
kg vs texts 638 1001 339 1978
GLM-4-9b-chat-hf
infobox vs texts 446 1352 46 1844
tables vs texts 728 1206 32 1966
kg vs texts 559 1356 46 1961
Mistral-7B-Instruct-v0.3
infobox vs texts 244 1675 53 1972
tables vs texts 647 1247 85 1979
kg vs texts 562 1350 58 1970
Table 3: Response Counts for Format Preference between Text and Structured Formats
C.5 Correlation Between Attention Gap and Dual-Claim Responses
To better understand the internal mechanisms underlying format bias, we analyze how attention allocation relates to the model’s ability to acknowledge both inputs. Specifically, for each input pair, we compute the difference in total attention mass assigned to the two evidence segments. We report the number of Both responses and the total number of samples within each group.
This analysis allows us to assess whether more balanced attention is associated with higher DCR. Table 4 summarizes the results. The observed negative correlation supports our hypothesis that early-stage processing asymmetry contributes to the exclusion of one input during generation.
Format Pair Avg Diff Both Total
Qwen3-8B
infoboxes vs tables 0.0134 292 1903
infoboxes vs texts 0.3126 112 1874
infoboxes vs KGs 0.1912 97 1964
tables vs texts 0.3020 159 1870
tables vs KGs 0.1672 110 1971
texts vs KGs -0.1265 209 1943
Mistral-7B-Instruct-v0.3
infoboxes vs tables -0.0667 276 1971
infoboxes vs texts 0.2928 54 1923
infoboxes vs KGs 0.1667 83 2030
tables vs texts 0.3765 107 1919
tables vs KGs 0.2455 116 2034
texts vs KGs -0.1511 100 1999
Llama-3.1-8B-Instruct
infoboxes vs tables -0.0207 711 1858
infoboxes vs texts 0.3451 468 1799
infoboxes vs KGs 0.1236 510 1900
tables vs texts 0.3697 615 1821
tables vs KGs 0.1591 587 1920
texts vs KGs -0.2183 343 1885
Table 4: Attention Imbalance and Dual Coverage Statistics Across Format Pairs and Models
C.6 Effects of Attention-Based Intervention
We evaluate the effect of our attention reweighting intervention by comparing model responses before and after modification. Table 5 shows the response distributions prior to intervention, and Table 6 shows the corresponding results after applying balanced attention. These results support our finding that the presence of bias can be reduced through inference-time control, while direction of bias remain relatively stable.
Format Pair Pref-A Pref-B Both Total
Qwen3-8B
tables vs texts 507 1204 159 1870
KGs vs texts 1089 645 209 1943
infoboxes vs tables 824 787 292 1903
infoboxes vs texts 543 1219 112 1874
KGs vs tables 1552 309 110 1971
infoboxes vs KGs 286 1581 97 1964
Mistral-7B-Instruct-v0.3
tables vs texts 413 1322 88 1823
KGs vs texts 905 994 100 1999
infoboxes vs tables 613 1000 256 1869
infoboxes vs texts 321 1548 54 1923
KGs vs tables 1356 562 116 2034
infoboxes vs KGs 296 1651 83 2030
Llama-3.1-8B-Instruct
tables vs texts 411 787 623 1821
KGs vs texts 807 750 334 1891
infoboxes vs tables 623 470 767 1860
infoboxes vs texts 492 825 487 1804
KGs vs tables 890 471 570 1931
infoboxes vs KGs 478 922 512 1912
Table 5: Response Counts before Attention Re-balacing
Format Pair Pref-A Pref-B Both Total
Qwen3-8B
tables vs texts 419 1138 314 1871
KGs vs texts 1106 564 269 1939
infoboxes vs tables 784 673 449 1906
infoboxes vs texts 512 1190 169 1871
KGs vs tables 1509 244 224 1977
infoboxes vs KGs 291 1501 175 1967
Mistral-7B-Instruct-v0.3
tables vs texts 378 1438 114 1930
KGs vs texts 882 970 143 1995
infoboxes vs tables 681 1014 283 1978
infoboxes vs texts 259 1611 57 1927
KGs vs tables 1371 526 139 2036
infoboxes vs KGs 244 1697 84 2025
Llama-3.1-8B-Instruct
tables vs texts 364 774 751 1889
KGs vs texts 767 728 472 1967
infoboxes vs tables 588 370 977 1935
infoboxes vs texts 496 810 592 1898
KGs vs tables 858 352 812 2022
infoboxes vs KGs 447 930 634 2011
Table 6: Response Counts after Attention Re-balacing
C.7 Format Preference across Topical Domains
To examine whether format bias generalizes across different types of content, we analyze format preference patterns within seven distinct topical domains. For each domain, we compute the FPR between all six pairs of formats and visualize the results using heatmaps.
Figures 8 and 9 present domain-level results for FPR and DCR, respectively. Each FPR heatmap summarizes the directional preference across format combinations, while the corresponding DCR plots capture the extent to which models jointly acknowledge conflicting evidence. The consistent trends across domains indicate that format bias manifests as a stable and systematic inductive pattern rather than a domain-specific artifact.
Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
infoboxes vs KGs 736 934 218 1888
infoboxes vs tables 812 477 560 1849
infoboxes vs texts 521 1110 171 1802
KGs vs tables 829 813 255 1897
KGs vs texts 780 956 124 1860
tables vs texts 453 1135 225 1813
GPT-4o-mini
infoboxes vs KGs 309 1342 242 1893
infoboxes vs tables 648 537 666 1851
infoboxes vs texts 377 1215 192 1784
KGs vs tables 1097 362 462 1921
KGs vs texts 689 838 335 1862
tables vs texts 382 971 455 1808
Qwen3-8B-FP8
infoboxes vs KGs 279 1582 109 1970
infoboxes vs tables 853 748 309 1910
infoboxes vs texts 508 1235 122 1865
KGs vs tables 1520 326 128 1974
KGs vs texts 1066 682 196 1944
tables vs texts 475 1197 189 1861
Qwen3-14B-FP8
infoboxes vs KGs 345 1418 176 1939
infoboxes vs tables 814 664 405 1883
infoboxes vs texts 509 1134 222 1865
KGs vs tables 1348 328 254 1930
KGs vs texts 921 724 274 1919
tables vs texts 399 1154 292 1845
Qwen3-32B-FP8
infoboxes vs KGs 261 1544 143 1948
infoboxes vs tables 749 661 484 1894
infoboxes vs texts 433 1286 130 1849
KGs vs tables 1488 254 219 1961
KGs vs texts 953 731 243 1927
tables vs texts 411 1218 228 1857
Table 7: Part 1 of Raw Response Counts for All Format Pairs across All Models
Format Pair Pref-A Pref-B Both Total
Qwen3-30B-A3B-FP8
infoboxes vs KGs 240 1423 111 1774
infoboxes vs tables 589 799 329 $17$
infoboxes vs texts 405 1167 105 1677
KGs vs tables 1351 310 120 1781
KGs vs texts 883 684 197 1764
tables vs texts 387 1127 172 1686
Gemma-2-9b-it-FP8
infoboxes vs KGs 590 1085 308 1983
infoboxes vs tables 796 532 601 1929
infoboxes vs texts 601 1170 104 1875
KGs vs tables 1094 570 327 1991
KGs vs texts 780 1009 167 1956
tables vs texts 453 1223 200 1876
Gemma-2-27b-it-FP8
infoboxes vs KGs 446 1192 396 2034
infoboxes vs tables 635 427 918 1980
infoboxes vs texts 592 1048 277 1917
KGs vs tables 1131 428 485 2044
KGs vs texts 893 708 401 2002
tables vs texts 522 1016 387 1925
GLM-4-9b-chat-hf
infoboxes vs KGs 841 1103 83 2027
infoboxes vs tables 1089 796 83 1968
infoboxes vs texts 594 1289 40 1923
KGs vs tables $12$ 791 29 2032
KGs vs texts 816 1098 69 1983
tables vs texts 489 1370 53 1912
Mistral-7B-Instruct-v0.3
infoboxes vs KGs 296 1651 83 2030
infoboxes vs tables 634 1061 276 1971
infoboxes vs texts 321 1548 54 1923
KGs vs tables 1356 562 116 2034
KGs vs texts 905 994 100 1999
tables vs texts 429 1383 107 1919
Table 8: Part 2 of Raw Response Counts for All Format Pairs across All Models
Group Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
KGs 4 vs 8 375 514 966 1855
12 vs 8 485 420 956 1861
12 vs 4 541 358 960 1859
infoboxes 4 vs 8 402 646 792 1840
12 vs 8 622 488 725 1835
12 vs 4 759 369 718 1846
tables 4 vs 8 634 743 473 1850
12 vs 8 681 586 582 1849
12 vs 4 769 611 470 1850
GPT-4o-mini
KGs 4 vs 8 348 414 1118 1880
12 vs 8 394 389 1103 1886
12 vs 4 438 324 1118 1880
infoboxes 4 vs 8 367 514 1001 1882
12 vs 8 494 441 937 1872
12 vs 4 616 375 887 1878
tables 4 vs 8 409 570 892 1871
12 vs 8 595 410 874 1879
12 vs 4 695 324 827 1846
Qwen3-8B-FP8
KGs 4 vs 8 382 499 1056 1937
12 vs 8 493 409 1034 1936
12 vs 4 559 319 1059 1937
infoboxes 4 vs 8 317 509 468 1294
12 vs 8 667 531 726 1924
12 vs 4 854 387 548 1789
tables 4 vs 8 533 846 550 1929
12 vs 8 820 528 581 1929
12 vs 4 1007 426 493 1926
Qwen3-14B-FP8
KGs 4 vs 8 361 517 1052 1930
12 vs 8 488 432 1013 1933
12 vs 4 536 359 1035 1930
infoboxes 4 vs 8 470 787 673 1930
12 vs 8 701 515 702 1918
12 vs 4 930 415 579 1924
tables 4 vs 8 497 773 658 1928
12 vs 8 716 493 716 1925
12 vs 4 951 396 581 1928
Qwen3-32B-FP8
KGs 4 vs 8 372 519 1023 1914
12 vs 8 459 417 901 1777
12 vs 4 548 340 1024 1912
infoboxes 4 vs 8 383 713 817 1913
12 vs 8 675 462 767 1904
12 vs 4 856 367 686 1909
tables 4 vs 8 484 758 670 1912
12 vs 8 709 455 612 1776
12 vs 4 937 379 590 1906
Table 9: Part 1 of Response Counts for Information Richness under Homogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Qwen3-30B-A3B-FP8
KGs 4 vs 8 378 561 988 1927
12 vs 8 529 434 963 1926
12 vs 4 580 348 998 1926
infoboxes 4 vs 8 401 683 840 1924
12 vs 8 633 484 799 1916
12 vs 4 851 371 697 1919
tables 4 vs 8 525 825 574 1924
12 vs 8 819 537 566 1922
12 vs 4 1021 415 482 1918
Gemma-2-9b-it-FP8
KGs 4 vs 8 445 482 1017 1944
12 vs 8 468 444 986 1898
12 vs 4 473 433 1016 1922
infoboxes 4 vs 8 434 549 962 1945
12 vs 8 549 518 875 1942
12 vs 4 649 467 822 1938
tables 4 vs 8 552 734 658 1944
12 vs 8 716 549 678 1943
12 vs 4 846 489 603 1938
Gemma-2-27b-it-FP8
KGs 4 vs 8 256 337 872 1465
12 vs 8 394 393 1190 1977
12 vs 4 471 300 1183 1954
infoboxes 4 vs 8 323 413 1257 1993
12 vs 8 414 398 1171 1983
12 vs 4 532 351 1105 1988
tables 4 vs 8 436 622 931 1989
12 vs 8 622 463 901 1986
12 vs 4 742 425 822 1989
GLM-4-9b-chat-hf
KGs 4 vs 8 472 662 853 1987
12 vs 8 633 538 810 1981
12 vs 4 703 454 828 1985
infoboxes 4 vs 8 568 960 451 1979
12 vs 8 898 769 306 1973
12 vs 4 1085 571 327 1983
tables 4 vs 8 721 1027 211 1959
12 vs 8 1005 753 217 1975
12 vs 4 1164 653 165 1982
Mistral-7B-Instruct-v0.3
KGs 4 vs 8 360 627 1007 1994
12 vs 8 530 435 1027 1992
12 vs 4 675 340 973 1988
infoboxes 4 vs 8 368 715 908 1991
12 vs 8 612 441 931 1984
12 vs 4 826 327 831 1984
tables 4 vs 8 573 755 662 1990
12 vs 8 870 537 580 1987
12 vs 4 1029 422 537 1988
Table 10: Part 2 of Response Counts for Information Richness under Homogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
KGs 0.45 vs un 401 445 1011 1857
0.9 vs un 507 404 942 1853
0.45 vs 0.9 404 441 1009 1854
infoboxes 0.45 vs un 450 534 848 1832
0.9 vs un 416 580 844 1840
0.45 vs 0.9 470 441 935 1846
tables 0.45 vs un 521 550 787 1858
0.9 vs un 526 551 777 1854
0.45 vs 0.9 516 502 833 1851
GPT-4o-mini
KGs 0.45 vs un 374 369 1140 1883
0.9 vs un 397 392 1093 1882
0.45 vs 0.9 371 394 1117 1882
infoboxes 0.45 vs un 380 439 1059 1878
0.9 vs un 398 488 993 1879
0.45 vs 0.9 435 431 1015 1881
tables 0.45 vs un 394 408 1079 1881
0.9 vs un 391 398 1094 1883
0.45 vs 0.9 400 407 1040 1847
Qwen3-8B-FP8
KGs 0.45 vs un 482 504 917 1903
0.9 vs un 554 485 751 1790
0.45 vs 0.9 485 522 925 1932
infoboxes 0.45 vs un 567 767 589 1923
0.9 vs un 79 117 60 256
0.45 vs 0.9 741 667 521 1929
tables 0.45 vs un 517 709 703 1929
0.9 vs un 572 688 672 1932
0.45 vs 0.9 604 579 749 1932
Qwen3-14B-FP8
KGs 0.45 vs un 433 456 1043 1932
0.9 vs un 454 488 991 1933
0.45 vs 0.9 470 463 999 1932
infoboxes 0.45 vs un 602 654 666 1922
0.9 vs un 606 747 575 1928
0.45 vs 0.9 620 659 642 1921
tables 0.45 vs un 432 560 937 1929
0.9 vs un 451 560 922 1933
0.45 vs 0.9 525 482 926 1933
Qwen3-32B-FP8
KGs 0.45 vs un 453 439 1022 1914
0.9 vs un 478 430 1008 1916
0.45 vs 0.9 437 465 1013 1915
infoboxes 0.45 vs un 475 534 902 1911
0.9 vs un 525 566 820 1911
0.45 vs 0.9 518 495 885 1898
tables 0.45 vs un 507 551 851 1909
0.9 vs un 465 509 787 1761
0.45 vs 0.9 530 531 853 1914
Table 11: Part 1 of Response Counts for Structure Quality under Homogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Qwen3-30B-A3B-FP8
KGs 0.45 vs un 482 472 974 1928
0.9 vs un 507 484 934 1925
0.45 vs 0.9 446 492 991 1929
infoboxes 0.45 vs un 477 570 874 1921
0.9 vs un 549 612 759 1920
0.45 vs 0.9 550 563 807 1920
tables 0.45 vs un 572 616 743 1931
0.9 vs un 569 634 725 1928
0.45 vs 0.9 564 555 807 1926
Gemma-2-9b-it-FP8
KGs 0.45 vs un 463 455 1020 1938
0.9 vs un 494 464 987 1945
0.45 vs 0.9 458 474 1012 1944
infoboxes 0.45 vs un 470 537 941 1948
0.9 vs un 487 639 817 1943
0.45 vs 0.9 531 529 886 1946
tables 0.45 vs un 575 583 746 1904
0.9 vs un 562 642 741 1945
0.45 vs 0.9 594 575 774 1943
Gemma-2-27b-it-FP8
KGs 0.45 vs un 351 409 1232 1992
0.9 vs un 374 415 1198 1987
0.45 vs 0.9 393 381 1217 1991
infoboxes 0.45 vs un 348 380 1266 1994
0.9 vs un 351 395 1246 1992
0.45 vs 0.9 372 362 1260 1994
tables 0.45 vs un 465 538 985 1988
0.9 vs un 479 559 949 1987
0.45 vs 0.9 514 500 977 1991
GLM-4-9b-chat-hf
KGs 0.45 vs un 565 530 889 1984
0.9 vs un 620 519 846 1985
0.45 vs 0.9 476 533 823 1832
infoboxes 0.45 vs un 558 748 423 1729
0.9 vs un 666 973 349 1988
0.45 vs 0.9 811 699 469 1979
tables 0.45 vs un 809 947 228 1984
0.9 vs un 774 994 211 1979
0.45 vs 0.9 842 820 318 1980
Mistral-7B-Instruct-v0.3
KGs 0.45 vs un 532 483 972 1987
0.9 vs un 635 518 811 1964
0.45 vs 0.9 545 558 873 1976
infoboxes 0.45 vs un 467 554 972 1993
0.9 vs un 474 679 820 1973
0.45 vs 0.9 585 508 900 1993
tables 0.45 vs un 559 642 791 1992
0.9 vs un 583 651 759 1993
0.45 vs 0.9 542 580 869 1991
Table 12: Part 2 of Response Counts for Structure Quality under Homogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
KGs 4 vs texts 556 1222 66 1844
8 vs texts 675 1085 85 1845
12 vs texts 672 858 104 1634
infoboxes 4 vs texts 482 1180 185 1847
8 vs texts 486 1160 193 1839
12 vs texts 546 1079 220 1845
tables 4 vs texts 426 1210 212 1848
8 vs texts 476 1117 251 1844
12 vs texts 532 1024 282 1838
GPT-4o-mini
KGs 4 vs texts 490 1076 298 1864
8 vs texts 577 894 361 1832
12 vs texts 645 779 437 1861
infoboxes 4 vs texts 323 1325 222 1870
8 vs texts 325 977 260 1562
12 vs texts 324 795 274 1393
tables 4 vs texts 324 1054 433 1811
8 vs texts 428 948 488 1864
12 vs texts 503 789 572 1864
Qwen3-8B-FP8
KGs 4 vs texts 693 1060 165 1918
8 vs texts 853 857 208 1918
12 vs texts 523 369 120 1012
infoboxes 4 vs texts 464 1352 100 1916
8 vs texts 529 1250 138 1917
12 vs texts 605 1169 140 1914
tables 4 vs texts 420 1341 164 1925
8 vs texts 443 1075 182 1700
12 vs texts 465 640 181 1286
Qwen3-14B-FP8
KGs 4 vs texts 591 1091 238 1920
8 vs texts 734 902 281 1917
12 vs texts 855 796 265 1916
infoboxes 4 vs texts 432 1288 198 1918
8 vs texts 525 1163 229 1917
12 vs texts 546 1114 253 1913
tables 4 vs texts 346 1279 294 1919
8 vs texts 429 1177 316 1922
12 vs texts 543 956 421 1920
Qwen3-32B-FP8
KGs 4 vs texts 673 978 244 1895
8 vs texts 764 875 270 1909
12 vs texts 807 713 238 1758
infoboxes 4 vs texts 384 1417 107 1908
8 vs texts 475 1281 145 1901
12 vs texts 516 1227 163 1906
tables 4 vs texts 369 1315 220 1904
8 vs texts 484 1178 244 1906
12 vs texts 600 992 314 1906
Table 13: Part 1 of Response Counts for Information Richness under Heterogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Qwen3-30B-A3B-FP8
KGs 4 vs texts 651 1108 159 1918
8 vs texts 807 888 219 1914
12 vs texts 841 685 198 1724
infoboxes 4 vs texts 392 1400 101 1893
8 vs texts 452 1310 145 1907
12 vs texts 515 1215 180 1910
tables 4 vs texts 434 1308 176 1918
8 vs texts 526 1158 229 1913
12 vs texts 684 947 281 1912
Gemma-2-9b-it-FP8
KGs 4 vs texts 488 1335 113 1936
8 vs texts 563 1227 141 1931
12 vs texts 626 1163 142 1931
infoboxes 4 vs texts 465 1362 105 1932
8 vs texts 520 1297 123 1940
12 vs texts 559 1249 125 1933
tables 4 vs texts 393 1392 152 1937
8 vs texts 464 1306 161 1931
12 vs texts 478 1230 226 1934
Gemma-2-27b-it-FP8
KGs 4 vs texts 571 1130 275 1976
8 vs texts 687 982 304 1973
12 vs texts 794 836 347 1977
infoboxes 4 vs texts 470 1351 162 1983
8 vs texts 542 1197 242 1981
12 vs texts 653 961 351 1965
tables 4 vs texts 506 1095 384 1985
8 vs texts 515 1040 435 1990
12 vs texts 603 822 481 1906
GLM-4-9b-chat-hf
KGs 4 vs texts 559 1371 38 1968
8 vs texts 616 1286 54 1956
12 vs texts 663 1209 78 1950
infoboxes 4 vs texts 444 1462 63 1969
8 vs texts 523 1397 48 1968
12 vs texts 600 1323 47 1970
tables 4 vs texts 433 1490 46 1969
8 vs texts 547 1368 57 1972
12 vs texts 674 1253 46 1973
Mistral-7B-Instruct-v0.3
KGs 4 vs texts 565 1322 69 1956
8 vs texts 764 1132 76 1972
12 vs texts 845 1015 103 1963
infoboxes 4 vs texts 233 1673 69 1975
8 vs texts 316 1608 52 1976
12 vs texts 391 1519 68 1978
tables 4 vs texts 373 1502 106 1981
8 vs texts 436 1452 84 1972
12 vs texts 551 1318 100 1969
Table 14: Part 2 of Response Counts for Information Richness under Heterogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Llama-3.1-8B-Instruct
KGs texts vs un 1207 556 74 1837
0.45 vs texts 535 1187 119 1841
0.9 vs texts 574 1158 112 1844
infoboxes texts vs un 1174 494 175 1843
0.45 vs texts 446 1216 185 1847
0.9 vs texts 394 1248 196 1838
tables texts vs un 910 722 216 1848
0.45 vs texts 679 928 237 1844
0.9 vs texts 646 905 295 1846
GPT-4o-mini
KGs texts vs un 1039 459 261 1759
0.45 vs texts 397 1177 281 1855
0.9 vs texts 410 1139 315 1864
infoboxes texts vs un 1334 328 200 1862
0.45 vs texts 297 1352 220 1869
0.9 vs texts 294 1349 220 1863
tables texts vs un 847 529 483 1859
0.45 vs texts 477 890 498 1865
0.9 vs texts 461 899 511 1871
Qwen3-8B-FP8
KGs texts vs un 1078 682 164 1924
0.45 vs texts 627 1139 159 1925
0.9 vs texts 612 1135 179 1926
infoboxes texts vs un 1282 423 92 1797
0.45 vs texts 320 1231 84 1635
0.9 vs texts 350 1469 99 1918
tables texts vs un 1051 683 196 1930
0.45 vs texts 588 1019 185 1792
0.9 vs texts 579 1140 208 1927
Qwen3-14B-FP8
KGs texts vs un 1090 601 225 1916
0.45 vs texts 582 1102 232 1916
0.9 vs texts 555 1112 251 1918
infoboxes texts vs un 1187 424 182 1793
0.45 vs texts 413 1320 181 1914
0.9 vs texts 383 1318 213 1914
tables texts vs un 961 582 373 1916
0.45 vs texts 539 1014 368 1921
0.9 vs texts 489 948 345 1782
Qwen3-32B-FP8
KGs texts vs un 992 683 233 1908
0.45 vs texts 619 1058 225 1902
0.9 vs texts 616 1047 248 1911
infoboxes texts vs un 1407 381 119 1907
0.45 vs texts 372 1424 113 1909
0.9 vs texts 366 1418 125 1909
tables texts vs un 904 711 292 1907
0.45 vs texts 644 988 275 1907
0.9 vs texts 609 1023 273 1905
Table 15: Part 1 of Response Counts for Structure Quality under Heterogeneous Format Conditions
Group Format Pair Pref-A Pref-B Both Total
Qwen3-30B-A3B-FP8
KGs texts vs un 1103 659 158 1920
0.45 vs texts 619 1106 188 1913
0.9 vs texts 664 1032 216 1912
infoboxes texts vs un 1449 378 85 1912
0.45 vs texts 360 1473 83 1916
0.9 vs texts 331 1484 101 1916
tables texts vs un 995 682 237 1914
0.45 vs texts 669 1008 236 1913
0.9 vs texts 643 $10$ 263 1916
Gemma-2-9b-it-FP8
KGs texts vs un 1326 487 125 1938
0.45 vs texts 448 1336 153 1937
0.9 vs texts 450 1347 137 1934
infoboxes texts vs un 1336 477 121 1934
0.45 vs texts 466 1379 90 1935
0.9 vs texts 467 1361 106 1934
tables texts vs un 1116 668 156 1940
0.45 vs texts 565 1199 171 1935
0.9 vs texts 542 1248 147 1937
Gemma-2-27b-it-FP8
KGs texts vs un 966 666 350 1982
0.45 vs texts 576 1032 376 1984
0.9 vs texts 496 1175 318 1989
infoboxes texts vs un 1338 477 165 1980
0.45 vs texts 418 1388 173 1979
0.9 vs texts 388 1413 186 1987
tables texts vs un 841 785 360 1986
0.45 vs texts 662 938 381 1981
0.9 vs texts 625 959 395 1979
GLM-4-9b-chat-hf
KGs texts vs un 1379 545 43 1967
0.45 vs texts 552 1378 40 1970
0.9 vs texts 604 1317 48 1969
infoboxes texts vs un 1497 422 47 1966
0.45 vs texts 432 1485 52 1969
0.9 vs texts 446 1473 53 1972
tables texts vs un 1229 700 38 1967
0.45 vs texts 637 1296 36 1969
0.9 vs texts 646 1280 40 1966
Mistral-7B-Instruct-v0.3
KGs texts vs un 1343 557 66 1966
0.45 vs texts 491 1419 58 1968
0.9 vs texts 477 1446 53 1976
infoboxes texts vs un 1692 245 42 1979
0.45 vs texts 260 1667 47 1974
0.9 vs texts 269 1656 54 1979
tables texts vs un 1244 654 80 1978
0.45 vs texts 645 1250 84 1979
0.9 vs texts 646 1231 100 1977
Table 16: Part 2 of Response Counts for Structure Quality under Heterogeneous Format Conditions
(a) Academia & Institutions
(b) Health & Medicine
(c) Social Sciences & Humanities
(d) Arts & Entertainment
(e) Occupations & Industry
(f) Government & Military
(g) Natural Sciences & Engineering
Figure 8: Domain-level FPR across format combinations.
(a) Academia & Institutions
(b) Health & Medicine
(c) Social Sciences & Humanities
(d) Arts & Entertainment
(e) Occupations & Industry
(f) Government & Military
(g) Natural Sciences & Engineering
Figure 9: Domain-level DCR across format combinations.